answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How many queen in a deck of 52 cards?

There are 4 queens are in a deck of 52 cards.

What is a continuous random variable?

In the simplest setting, a continuous random variable is one that can assume any value on some interval of the real numbers.

For example, a uniform random variable is often defined on the unit interval [0,1], which means that this random variable could assume any value between 0 and 1, including 0 and 1. Some possibilities would be 1/3, 0.3214, pi/4, e/5, and so on ... in other words, any of the numbers in that interval.

As another example, a normal random variable can assume any value between -infinity and +infinity (another interval). Most of these values would be extremely unlikely to occur but they would be possible. The random variable could assume values of 3, -10000, pi, 1000*pi, e*e, ... any possible value in the real numbers.

It is also possible to define continue random variables that assume values on the entire (x,y) plane, or just on the circumference of a circle, or anywhere that you can imagine that is essentially equivalent (in some sense) to pieces of a real line.

How many favorable outcomes would be possible if the number of total possible outcomes is increased to 24?

The answer depends on what the trials are, what constitutes a favourable outcome, and how many possible outcomes there were previously. And since you have not bothered to provide these crucial bits of information, I cannot provide a more useful answer.

The answer depends on what the trials are, what constitutes a favourable outcome, and how many possible outcomes there were previously. And since you have not bothered to provide these crucial bits of information, I cannot provide a more useful answer.

The answer depends on what the trials are, what constitutes a favourable outcome, and how many possible outcomes there were previously. And since you have not bothered to provide these crucial bits of information, I cannot provide a more useful answer.

The answer depends on what the trials are, what constitutes a favourable outcome, and how many possible outcomes there were previously. And since you have not bothered to provide these crucial bits of information, I cannot provide a more useful answer.

Why is calculating the standard deviation an important part of data analysis?

It is a measure of the spread of the results around their expected value.

It is a measure of the spread of the results around their expected value.

It is a measure of the spread of the results around their expected value.

It is a measure of the spread of the results around their expected value.

What proportion of a normal distribution is located between z-0.90 and z 0.90?

0.368 or 36.8%.

And you should specify that it is a standard normal distribution.

0.368 or 36.8%.

And you should specify that it is a standard normal distribution.

0.368 or 36.8%.

And you should specify that it is a standard normal distribution.

0.368 or 36.8%.

And you should specify that it is a standard normal distribution.

What is the probability of not geeting a queen in a well shuffled deck of 52 playing cards?

If only one card is selected the probability is 12/13.

If only one card is selected the probability is 12/13.

If only one card is selected the probability is 12/13.

If only one card is selected the probability is 12/13.

Why does the standard deviation get smaller as the individual in a group score more similarly on a test?

Because the standard deviation is a measure of the spread in scores. As individuals score more similarly, the spread gets smaller.



Because the standard deviation is a measure of the spread in scores. As individuals score more similarly, the spread gets smaller.



Because the standard deviation is a measure of the spread in scores. As individuals score more similarly, the spread gets smaller.



Because the standard deviation is a measure of the spread in scores. As individuals score more similarly, the spread gets smaller.

What are the characteristics of probability?

Probability is a numerical value defined for a set of outcomes. It is non-negative and such that the sum or integral over all possible outcomes is 1.

How does the experimental result differ from the theoretical in terms of accuracy?

Provided that the correct model is used, the theoretical probability is correct. The experimental probability tends towards the theoretical value as the number of trials increases.

Provided that the correct model is used, the theoretical probability is correct. The experimental probability tends towards the theoretical value as the number of trials increases.

Provided that the correct model is used, the theoretical probability is correct. The experimental probability tends towards the theoretical value as the number of trials increases.

Provided that the correct model is used, the theoretical probability is correct. The experimental probability tends towards the theoretical value as the number of trials increases.

What is the probability that the a dice 2 is drawn if the two dice total is 5?

The probability, when the 2-dice total is 5, that one of the two dice shows a two is 1/2. The probability that that die is selected is 1/4.

The probability, when the 2-dice total is 5, that one of the two dice shows a two is 1/2. The probability that that die is selected is 1/4.

The probability, when the 2-dice total is 5, that one of the two dice shows a two is 1/2. The probability that that die is selected is 1/4.

The probability, when the 2-dice total is 5, that one of the two dice shows a two is 1/2. The probability that that die is selected is 1/4.

What is the probability of flipping heads and rolling an odd number?

These would be independent events; therefore, we can multiply the probabilities of each of the two events.

Probability of flipping a head: 1/2

Probability of rolling an odd number with a single die: 1/6

Required probability : 1/2 x 1/6 = 1/12

How many different committees can be formed from 10 teachers and 30 students if the committee consists of 2 teachers and 2 students?

There are (10 x 9)/2 = 45 different possible pairs of 2 teachers. For each of these . . .

There are (30 x 29)/2 = 435 different possible pairs of students.

The total number of different committees that can be formed is (45 x 435) = 19,575 .

How do you see that the sampling distribution of the mean is normal?

According to the Central Limit Theorem, the mean of a sufficiently large number of independent random variables which have a well defined mean and a well defined variance, is approximately normally distributed.


The necessary requirements are shown in bold.



According to the Central Limit Theorem, the mean of a sufficiently large number of independent random variables which have a well defined mean and a well defined variance, is approximately normally distributed.


The necessary requirements are shown in bold.



According to the Central Limit Theorem, the mean of a sufficiently large number of independent random variables which have a well defined mean and a well defined variance, is approximately normally distributed.


The necessary requirements are shown in bold.



According to the Central Limit Theorem, the mean of a sufficiently large number of independent random variables which have a well defined mean and a well defined variance, is approximately normally distributed.


The necessary requirements are shown in bold.

What is the probability of getting an ace of hearts and a king of hearts from a regular deck of cards without jokers?

The answer depends on how many cards are drawn and whether or not they are replaced before the next card is drawn.

The answer depends on how many cards are drawn and whether or not they are replaced before the next card is drawn.

The answer depends on how many cards are drawn and whether or not they are replaced before the next card is drawn.

The answer depends on how many cards are drawn and whether or not they are replaced before the next card is drawn.

What is the probability of getting 4 red cards from a standard deck of 52 cards?

If you are picking a card and putting it back before picking the next, it would be 1/2 * 1/2 * 1/2 * 1/2 = 1/16

If you are not replacing the cards, it would be 26/52 * 25/51 * 24/50 * 23/49 = 358800 / 6497400, or approximately 5.5%.

How many ways word crisis arranged?

There are 6 letter in the word crisis so the answer seems to be 6! (6 factorial), but there are two 'i's and two 's's which are indistinguishable and each of which can be arranged in 2! ways. So:

number_of_ways = 6!/2!2!
= 180 ways.

How many different ways can you arrange the letters of the word Malayalam?

There are 9!/[2!*4!*2!] = 3780 if you treat M and m as equivalent. Otherwise, there are twice as many.

What is probability compliment?

"That was a good bet!", perhaps.

"That was a good bet!", perhaps.

"That was a good bet!", perhaps.

"That was a good bet!", perhaps.

What is the values of the skewdness and kurtosis coefficient for the normal distribution 0 and 3 respectively?

No.


Skewness is 0, but kurtosis is -3, not 3.



No.


Skewness is 0, but kurtosis is -3, not 3.



No.


Skewness is 0, but kurtosis is -3, not 3.



No.


Skewness is 0, but kurtosis is -3, not 3.

How many possible outcomes for normal distribution?

Infinitely many. The normal distribution is applicable to a continuous variable whose domain is the whole of the real numbers.



Infinitely many. The normal distribution is applicable to a continuous variable whose domain is the whole of the real numbers.



Infinitely many. The normal distribution is applicable to a continuous variable whose domain is the whole of the real numbers.



Infinitely many. The normal distribution is applicable to a continuous variable whose domain is the whole of the real numbers.